11

The marginal labour cost function is given by the equation

MLC = 3 + 4L.

Calculate the cost of employing the first eight labourers.

a.       −114

b.       114

c.       152

d.       119

 

12

If the demand function is

P = Q2 − 7Q − 8,

where P and Q are the price and quantity respectively, determine the point price elasticity of demand (rounded to three decimal places) if P = 51. Is demand elastic or inelastic at this price?

[a.   ed = −0.075;inelastic [b.  ed = −0.255;inelastic

[c.   ed = 1.300; elastic

[d.  ed = −0.253; inelastic 


1
Expert's answer
2021-06-15T09:01:35-0400

Solution.

1.


"MLC=3+4L"

"C=\int\limits_0^8 (3+4L)dL=\newline 3L+2L^2|_0^8=3•8+2•64=24+128=152."

Answer. c.152

2.

"P=Q^2-7Q-8"

If P=51, then "Q^2-7Q-59=0."

From here "Q=-4.941," or "Q=11.941."

Find "e_d" for formula:


"e_d=\frac{dQ}{dP}\frac{P}{Q}"

Find P(Q):

"Q^2-7Q-8-P=0"

"D=49-4(-8-P)=81+4P"

"Q=\frac{7\pm\sqrt{81+4P}}{2}"

"e_d=\pm\frac{1}{\sqrt{81+4P}}\frac{51}{11.941}=\pm 0.253"

"|e_d|<1," then demand inelastic.

Answer. d.-0.253, inelastic.